# Simulation of depth-dose curves and water equivalent ratios of energetic proton beams in cortical bone

### v1

@@ Abstract
We have determined the depth-dose curve, the penetration range, and the water equivalent ratio (WER), for proton beams of clinical energies in cortical bone, by means of a detailed and accurate simulation that combines molecular dynamics and Monte Carlo techniques. The fundamental input quantities (stopping power and energy loss straggling) for the simulation were obtained from a reliable electronic excitation spectrum of the condensed-phase target, which takes into account the organic and mineral phases that form it. Our simulations with these inputs, that are in excellent agreement with the scarce data available for a cortical bone target, deviate from simulations performed using other stopping quantities, such as those provided in the widely used ICRU Report 49. The results of this work emphasize the importance of an accurate determination of the stopping quantities of cortical bone in order to advance towards the millimetric precision for the proton penetration ranges and deposited dose needed in radiotherapy.
  • April 25, 2024
@@ 1 Introduction
The aim of the present work is to study the main quantities related to the energy deposition by protons, having clinical energies, in both liquid water and cortical bone. From Monte Carlo simulations with the SEICS code [23, 24], we get the depth-dose curves, the penetration range and the WER in cortical bone, in order to compare them with the corresponding fot liquid water. To account for the condensed-phase nature of the studied materials, we use the MELF-GOS method (based on the dielectric formalism) to calculate the ELF of the target as well as a multi-phase approach for cortical bone [4]. This leads to stopping power values showing differences of the order of 1010101015151515% between our calculations (based in the dielectric formalism and the MELF-GOS methodology) and ICRU tabulation [21, 22] below 200 keV for water and 1000 keV for bone. Despite the scarcity of experimental data of depth-dose curves in bone, our depth-dose simulation is in excellent agreement with Vasiliev et al.’s experimental measurements [25] , even better than using ICRU data. In addition, our simulations of WER in cortical bone show small differences between using the MELF-GOS stopping power or the ICRU one. In spite of a fair agreement with recently reported WER values by Burin et al. [8], the differences between both datasets are more visible.
@@ 3 Results and discussion
Figure 5: Depth-dose curve in cortical bone, for an initial $200$ MeV proton beam, previously passed through a $100$ mm-wide target of liquid water. Symbols depict the results of experimental measurements in a cortical bone phantom [25], while red solid (black dashed) line corresponds to a SEICS simulation using the MELF-GOS (ICRU) stopping power. All curves are normalized to unit area. The inset sketches the experimental setup, in which a $200$ MeV ± $\pm\,0.3$ initial proton beam traverses a liquid water target, resulting in its energy spectrum widened before entering the cortical bone target.
@@
The experimental information both on the stopping power and on the depth-dose curves for proton beams in cortical bone is very scarce, despite the importance of this target in radiotherapy treatments. The only measured depth-dose curve we are aware of is that of $200$ MeV protons interacting with a cortical bone phantom material reported by Vasiliev et al. [25]. In the experimental setup, the $200$ MeV beam, having an energy width of ±$\pm 0.3$% , was first passed through a liquid water energy degrader $100$ mm in depth. The resulting, further energy widened, beam then bombarded the cortical bone phantom. This setup is sketched in the inset of Fig. 5.
@@
The SEICS code allows to introduce an energy spectrum to the incident beam, as well as to measure the energy spectrum of the protons leaving the target slab. The spectra of the protons entering and leaving the liquid water target of $100$ mm width are depicted in the inset of Fig. 5. The energy loss straggling of liquid water makes the beam to increase its energy width from the initial ±$\pm 0.3$% to ±$\pm 0.93$% when exiting the liquid water slab. This latter value has been then used to simulate the proton propagation in the cortical bone target. The red solid line in Fig. 5 corresponds to the simulation using the stopping powers and energy loss stragglings obtained in this work (both for liquid water and cortical bone), while the black dashed line uses the ICRU stopping powers, complemented with the current energy loss straggling values. As can be clearly seen, MELF-GOS slowing down quantities yield a simulated depth-dose curve which almost perfectly agrees with the experimentally measured data from Ref. [25], except for a slight overestimation of the dose around the Bragg peak maximum (again, results have been normalized, so as the areas under the depth-dose curves have a value of one). Such a good result (particularly, the excellent reproduction of the depth for maximum dose) has been obtained without the need of slight adjustments in the nominal initial beam energy, which sometimes are needed due to the initial energy uncertainties [5]. In contrast, the simulation using the ICRU stopping power (for both target materials) results in a Bragg curve maximum slightly shifted to shallower depths, together with a somewhat more intense dose around the maximum, as easily observed in the figure. The range variation associated with the use of one set of stopping powers or the other is of $1.3$ mm. This comparison supports the accuracy of the slowing down quantities calculated by means of the MELF-GOS method, as well as the good performance of the SEICS simulation code.

### v2

@@ Abstract
We have determined the depth-dose curve, the penetration range, and the water equivalent ratio (WER), for proton beams of clinical energies in cortical bone, by means of a detailed and accurate simulation that combines molecular dynamics and Monte Carlo techniques. The fundamental input quantities (stopping power and energy loss straggling) for the simulation were obtained from a reliable electronic excitation spectrum of the condensed-phase target, which takes into account the organic and mineral phases that form it. Our simulations with these inputs, that are in excellent agreement with the scarce data available for a cortical bone target, deviate from simulations performed using other stopping quantities, such as those provided by the International Commission on Radiation Units and Measurements (ICRU) in its widely used Report 49. The results of this work emphasize the importance of an accurate determination of the stopping quantities of cortical bone in order to advance towards the millimetric precision for the proton penetration ranges and deposited dose needed in radiotherapy.
  • September 17, 2024
@@ 1 Introduction
The aim of the present work is to study the main quantities related to the energy deposition by protons, having clinical energies, in both liquid water and cortical bone. From Monte Carlo simulations with the Simulation of Energetic Ions and Clusters Through Solids (SEICS) code [23, 24], we get the depth-dose curves, the penetration range and the WER in cortical bone, in order to compare them with the corresponding quantities for liquid water. To account for the condensed-phase nature of the studied materials, we use the Mermin Energy Loss Function – Generalised Oscillator Strengths (MELF-GOS) method [25, 26] (based on the dielectric formalism) to calculate the electronic excitation spectrum of the target (encoded in its Energy Loss Function, ELF) as well as a multi-phase approach for cortical bone [4]. This leads to stopping power values showing differences of the order of 1010101015151515% between our calculations (based in the dielectric formalism and the MELF-GOS methodology) and tabulation by the International Commission on Radiation Units and Measurements (ICRU) [21, 22] below 200 keV for water and 1000 keV for bone. Despite the scarcity of experimental data of depth-dose curves in bone, our depth-dose simulation is in excellent agreement with Vasiliev et al.’s experimental measurements [27] , even better than using ICRU data. In addition, our simulations of WER in cortical bone show small differences between using the MELF-GOS stopping power or the ICRU one. In spite of a fair agreement with recently reported WER values by Burin et al. [8], the differences between both datasets are more visible.
@@ 3 Results and discussion
Figure 5: (a) Depth-dose curve in cortical bone, for an initial $200$ MeV proton beam, previously passed through a $100$ mm-wide target of liquid water. Symbols depict the results of experimental measurements in a cortical bone phantom [27], while red solid (black dashed) line corresponds to a SEICS simulation using the MELF-GOS (ICRU) stopping power. All curves are normalized to unit area. The inset sketches the experimental setup, in which a $200$ MeV ± $\pm\,0.3$ initial proton beam traverses a liquid water target, resulting in its energy spectrum widened before entering the cortical bone target. (b) Detail of the depth-dose curves around the maximum dose, where arrows indicate the location of the penetration range for each curve, estimated as the depth at which the dose falls to $80$ after the maximum.
@@
The experimental information both on the stopping power and on the depth-dose curves for proton beams in cortical bone is very scarce, despite the importance of this target in radiotherapy treatments. The only measured depth-dose curve we are aware of is that of $200$ MeV protons interacting with a cortical bone phantom material reported by Vasiliev et al. [27]. In the experimental setup, the $200$ MeV beam, having an energy width of ±$\pm 0.3$% , was first passed through a liquid water energy degrader $100$ mm in depth. The resulting, further energy widened, beam then bombarded the cortical bone phantom. This setup is sketched in the inset of Fig. 5(a).
@@
The SEICS code allows to introduce an energy spectrum to the incident beam, as well as to measure the energy spectrum of the protons leaving the target slab. The spectra of the protons entering and leaving the liquid water target of $100$ mm width are depicted in the inset of Fig. 5(a). The energy loss straggling of liquid water makes the beam to increase its energy width from the initial ±$\pm 0.3$% to ±$\pm 0.93$% when exiting the liquid water slab. This latter value has been then used to simulate the proton propagation in the cortical bone target. The red solid line in Fig. 5(a) corresponds to the simulation using the stopping powers and energy loss stragglings obtained in this work (both for liquid water and cortical bone), while the black dashed line uses the ICRU stopping powers, complemented with the current energy loss straggling values. As can be clearly seen, MELF-GOS slowing down quantities yield a simulated depth-dose curve which almost perfectly agrees with the experimentally measured data from Ref. [27], except for a slight overestimation of the dose around the Bragg peak maximum (again, results have been normalized, so as the areas under the depth-dose curves have a value of one). Such a good result (particularly, the excellent reproduction of the depth for maximum dose) has been obtained without the need of slight adjustments in the nominal initial beam energy, which sometimes are needed due to the initial energy uncertainties [5]. In contrast, the simulation using the ICRU stopping power (for both target materials) results in a Bragg curve maximum slightly shifted to shallower depths, together with a somewhat more intense dose around the maximum, as easily observed in the figure. Figure 5(b) focuses around the maximum of the depth-dose curve, in order to better appreciate the scale of the range differences. The range obtained from each curve has been marked by arrows. Clearly, the range obtained using current stopping quantities ($\sim 9.71$ cm) is in closer agreement to the experimental value ($\sim 9.75$ cm) than that obtained from simulations using ICRU stopping power ($\sim 9.58$ cm). The range variation associated with the use of one set of stopping powers or the other is of $\sim 1.3$ mm. This comparison supports the accuracy of the slowing down quantities calculated by means of the MELF-GOS method, as well as the good performance of the SEICS simulation code.
@@ 4 Summary and conclusions
The datasets generated during the current study and simulation code are available from the corresponding author on reasonable request.